OpenStudy (anonymous):

Evaluate the following limits. If needed, enter INF for \infty and MINF for -\infty. (a) \lim_{ x \rightarrow \infty } \frac {10 x + 9 }{ 11 x^2- 3 x + 4 } (b) \lim_{ x \rightarrow -\infty } \frac {10 x + 9 }{ 11 x^2- 3 x + 4 }

OpenStudy (slaaibak):

Is it just me, or is the latex not working?

OpenStudy (zarkon):

\[x^2\] working for me

OpenStudy (shayaan_mustafa):

latex is not workin @rukh

OpenStudy (slaaibak):

I mean rukh's post. Zarkon, yours is working for me.

OpenStudy (anonymous):

its been like this for a while now

OpenStudy (zarkon):

(a) \[\lim_{ x \rightarrow \infty } \frac {10 x + 9 }{ 11 x^2- 3 x + 4 }\] (b) \[\lim_{ x \rightarrow -\infty } \frac {10 x + 9 }{ 11 x^2- 3 x + 4 }\]

OpenStudy (anonymous):

start with "\[" end with "\]"

OpenStudy (anonymous):

@Zarkon thanks

OpenStudy (anonymous):

\[\lim_{ x \rightarrow \infty } \frac {10 x + 9 }{ 11 x^2- 3 x + 4 }\]

OpenStudy (anonymous):

oops what zarkon said

OpenStudy (anonymous):

you have to start the latex with \[

OpenStudy (anonymous):

and end with \]

OpenStudy (anonymous):

then it should work

OpenStudy (zarkon):

or \(

OpenStudy (zarkon):

and \)

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

\(\sum\)

OpenStudy (anonymous):

who knew?

OpenStudy (zarkon):

yes...to do inline stuff if \(x^2\) is greater than \(x^3\)...

OpenStudy (zarkon):

doesn't put it on a new line

OpenStudy (anonymous):

i think the answer might be 0

OpenStudy (anonymous):

what on earth? if \(x^2\) is greater than \(x^3\)

OpenStudy (zarkon):

it is

OpenStudy (anonymous):

sorry yes, the answer is zero to both since degree of numerator is smaller than degree of denominator

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

i learn something new every day does require that extra "shift" though

OpenStudy (anonymous):

http://openstudy.com/study?signup#/updates/4f7761e7e4b0ddcbb89dced0 help with this one anyone?

OpenStudy (zarkon):

if you are doing inline latex you can also do the following ...since inline doesn't look as good at not inline \(\sum_{n=1}^{\infty}a_n\) vs \(\displaystyle\sum_{n=1}^{\infty}a_n\) \sum_{n=1}^{\infty}a_n vs \displaystyle\sum_{n=1}^{\infty}a_n

OpenStudy (anonymous):

i will definitely use it because there are times when it is annoying to have to start a new line

OpenStudy (anonymous):

i noticed that with actual latex documents using $ instead of \(

OpenStudy (zarkon):

yes...same thing also $$ for \[

OpenStudy (anonymous):

can you use \( instead of $ in a document?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

http://openstudy.com/study?signup#/updates/4f7761e7e4b0ddcbb89dced0 help with this one anyone?

OpenStudy (anonymous):

i think i remember reading that $ was actually a hack and the preferred method is \[ but i could be wrong

OpenStudy (zarkon):

$ is for \( $$ is for \[

OpenStudy (anonymous):

ah i see. all this time i have been using $ and \[ but never the other two thanks!

OpenStudy (zarkon):

I used $ for years...just started to use \( full time now

OpenStudy (anonymous):

more key strokes

OpenStudy (zarkon):

np :)

OpenStudy (zarkon):

I think it looks nicer though ;)

OpenStudy (anonymous):

i think i will post this in latex practicing